OpenStudy (anonymous):

fifth root of 8 multiplied by square root of 8 over cube root of 8 to the power of 5 8 to the power of negative 3 over 10 8 to the power of negative 41 over 30 8 to the power of negative 29 over 30 8 to the power of 71 over 30

OpenStudy (anonymous):

\[5\sqrt{8}*\sqrt{8}\over3\sqrt{8^5}\]

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (anonymous):

\[5\sqrt{2}/90\]

OpenStudy (anonymous):

96, not 90. Sorry.

OpenStudy (anonymous):

how would I solve that?

OpenStudy (campbell_st):

write each term in index form and apply the index laws for multiplication and division \[8^{\frac{1}{5}} \times 8^{\frac{1}{2}} \div 8^{\frac{1}{3}}\]

OpenStudy (anonymous):

So 8.37?

OpenStudy (campbell_st):

oops \[8^{\frac{1}{5}} \times 8^{\frac{1}{2}} \div 8 ^{\frac{5}{3}}\]

OpenStudy (campbell_st):

sorry about missing the power of 5...

OpenStudy (midhun.madhu1987):

 can u try to solve it further?